jose wants to use a glass rod as an insulator. the rods diameter is 9 mm and kts resistivity is 1×10^9m. how long must be the rod to he offer a resistance of 9.1×10^10 ohms?


1
Expert's answer
2021-03-26T11:28:00-0400

"R=\rho \frac{l}{A}\to l=\frac{RA}{\rho}"


"A=\pi d^2/4=3.14\cdot0.009^2/4=6.36\cdot10^{-5}\ (m^2)"


"l=\frac{RA}{\rho}=\frac{9.1\cdot10^{10}\cdot 6.36\cdot 10^{-5}}{1\cdot10^9}\approx0.006\ (m^2)=6\ (mm)" . Answer